If a 50g sample of substance Z decays to 47g in 11 days, what is the half-life of substance Z? Give your answer to the nearest day.
Answer by amoresroy(361)

half-life = elapsed time * log 2 / log (beginning amount/ending amount)

elapsed time = 11/365 = 0.030136986
beginning amount = 50g
ending amount = 47g
Use a scientific calculator to solve this.
half-life = 0.030136986 * log2 / log (50g/47g)
= 0.337603726
= 365*0.337603726
= 123.23 days
= 123 days